# Constraining Zero-Point Length from Gravitational Baryogenesis

### v1

@@ I Introduction
$\phi(r)=-\frac{M}{\sqrt{r^{2}+l_{0}^{2}}},$ (2)
where $M$ is the mass of the source of the potential and for simplicity, we have set $G=\hbar=c=1$. Thus, a point particle with mass $m$ at distance $r=R$ from mass $M$, feels the force
$\vec{F}=-m\vec{\nabla}\phi(r)|_{r=R}=-\frac{Mm}{R^{2}}\left[1+\frac{l_{0}^{2}}{R^{2}}\right]^{-3/2}\hat{r}$. (3)
@@
While direct experimental verification of such Planck-scale effects remains elusive, their potential implications for cosmology and early universe physics provide a fertile ground for testing these models. However, cosmological scenarios incorporating a fundamental zero-point length cannot be constrained using observational bounds from Big-Bang nucleosynthesis (BBN), in contrast to analyses such as Refs. Ava1 ; Ava2 ; Luciv . The reason is that the energy (and corresponding length and time) scales at which zero-point length effects become relevant lie far above the BBN scale, rendering primordial nucleosynthesis completely insensitive to such corrections. Consequently, in this work we employ gravitational baryogenesis which operates at energies many orders of magnitude higher than BBN as a viable probe for constraining zero-point length cosmology.
@@ II Modified cosmology through zero-point length correction
$S_{h}$ $=$ $\pi R^{2}\left(1+\frac{l_{0}^{2}}{R^{2}}\right)^{-1/2}+3\pi l_{0}^{2}\left(1+\frac{l_{0}^{2}}{R^{2}}\right)^{-1/2}$ (5)
$-3\pi l_{0}^{2}\ln\left(R+\sqrt{R^{2}+l_{0}^{2}}\right).$
@@
$dS_{h}=2\pi R\left(1+\frac{l_{0}^{2}}{R^{2}}\right)^{-3/2}dR.$ (6)
Although the associated entropy with the apparent horizon is complicated, for $l_{0}^{2}\rightarrow 0$ it recovers the area law of entropy, namely $S_{h}=\pi R^{2}$, as expected. It is evident that $l_{0}$ must carry the dimension of length (or inverse energy in natural units), since the corrected entropy-area expression must remain dimensionless in natural units. Note that the area of the apparent horizon is $A=4\pi R^{2}$, while the volume enveloped by the apparent horizon is given by $V=4\pi R^{3}/3$. Note that here $R$ is the apparent horizon radius defined as Hay1 ; Hay2 ; Bak
@@
$\frac{2}{R^{3}}\left(1+\frac{l_{0}^{2}}{R^{2}}\right)^{-3/2}dR=\frac{8\pi}{3}d\rho.$ (13)
@@
$\frac{2}{l_{0}^{2}}\left[1-\frac{l_{0}^{2}}{2R^{2}}+\frac{3}{8}\frac{l_{0}^{4}}{R^{4}}+...\right]+C=\frac{8\pi}{3}\rho,$ (14)
@@
$H^{2}+\frac{k}{a^{2}}-\alpha\left(H^{2}+\frac{k}{a^{2}}\right)^{2}=\frac{8\pi}{3}(\rho+\rho_{\Lambda}),$ (16)
where $\alpha=3l_{0}^{2}/4$, $\rho_{\Lambda}=\Lambda/8\pi$. In the absence of the cosmological constant ($\Lambda=0$) and for a flat universe ($k=0$), the modified Friedmann equation is simplified as
$H^{2}-\alpha H^{4}=\frac{8\pi}{3}\rho.$ (17)
@@
$\alpha X^{2}+X-\frac{8\pi}{3}\rho=0,$ (18)
where we have again set $\rho_{\Lambda}=0=\Lambda$, for simplicity. Eq. (18) admits the following solution
$X=\frac{2}{3l_{0}^{2}}\left[1\mp\sqrt{1-8\pi l_{0}^{2}\rho}\right].$ (19)
@@
$H^{2}+\frac{k}{a^{2}}=\frac{8\pi}{3}\rho\,(1+\Gamma\rho),$ (21)
where we have defined $\Gamma=2\pi l_{0}^{2}$.
@@
$H^{2}\approx\frac{8\pi}{3}\Gamma\rho^{2}.$ (22)
@@ III Baryogenesis Constraints on Zero-Point Length Cosmology
Beginning with the modified first Friedmann equation ($G\neq 1$)
@@ IV The relation between cosmic time and temperature in zero-point length cosmology
which is the time-temperature relation in zero-point length cosmology. In the limiting case where $l_{0}\to 0$ ($\alpha=0$), the time-temperature relation simplifies to $t=\frac{1}{2\beta T^{2}}$, which corresponds exactly to the standard general relativity result Weinberg . We note that while our modified Friedmann equations were originally derived in natural units, dimensional consistency necessitates that the right-hand side of Eq. (56) must possess dimensions of time $[t]$, and accordingly all fundamental constants have been properly restored.
@@
Figure 2: The behavior of the temperature vs time in the radiation dominated era of zero-point length cosmology for different values of the zero-point length parameter $\tilde{l}_{0}=[0.5\times 10^{-10},7\times 10^{-10},10^{-9}]$.

### v2

@@ I Introduction
$\phi(r)=-\frac{GM}{\sqrt{r^{2}+l_{0}^{2}}},$ (2)
where $M$ is the mass of the source of the potential and for simplicity, we have set $\hbar=c=1$. Thus, a point particle with mass $m$ at distance $r=R$ from mass $M$, feels the force
$\vec{F}=-m\vec{\nabla}\phi(r)|_{r=R}=-\frac{GMm}{R^{2}}\left[1+\frac{l_{0}^{2}}{R^{2}}\right]^{-3/2}\hat{r}$. (3)
@@
While direct experimental verification of such Planck-scale effects remains elusive, their potential implications for cosmology and early universe physics provide a fertile ground for testing these models. However, cosmological scenarios incorporating a fundamental zero-point length cannot be constrained using observational bounds from Big-Bang nucleosynthesis (BBN), in contrast to analysis such as Refs. Barrow2 ; Anish ; Ava1 ; Ava2 . The reason is that the energy (and corresponding length and time) scales at which zero-point length effects become relevant lie far above the BBN scale, rendering primordial nucleosynthesis completely insensitive to such corrections. Consequently, in this work we employ gravitational baryogenesis which operates at energies many orders of magnitude higher than BBN as a viable probe for constraining zero-point length cosmology Luciv ; Luciv2 ; Luciv3 .
@@ II Modified cosmology through zero-point length correction
$S_{h}$ $=$ $\frac{\pi R^{2}}{G}\left(1+\frac{l_{0}^{2}}{R^{2}}\right)^{-1/2}+\frac{3\pi l_{0}^{2}}{G}\left(1+\frac{l_{0}^{2}}{R^{2}}\right)^{-1/2}$ (5)
$-\frac{3\pi l_{0}^{2}}{G}\ln\left(R+\sqrt{R^{2}+l_{0}^{2}}\right).$
@@
$dS_{h}=\frac{2\pi R}{G}\left(1+\frac{l_{0}^{2}}{R^{2}}\right)^{-3/2}dR.$ (6)
Although the associated entropy with the apparent horizon is complicated, for $l_{0}^{2}\rightarrow 0$ it recovers the area law of entropy, namely $S_{h}=\pi R^{2}/G$, as expected. It is evident that $l_{0}$ must carry the dimension of length (or inverse energy in natural units), since the corrected entropy-area expression must remain dimensionless in natural units. Note that the area of the apparent horizon is $A=4\pi R^{2}$, while the volume enveloped by the apparent horizon is given by $V=4\pi R^{3}/3$. Note that here $R$ is the apparent horizon radius defined as Hay1 ; Hay2 ; Bak
@@
$\frac{2}{R^{3}}\left(1+\frac{l_{0}^{2}}{R^{2}}\right)^{-3/2}dR=\frac{8\pi G}{3}d\rho.$ (13)
@@
$\frac{2}{l_{0}^{2}}\left[1-\frac{l_{0}^{2}}{2R^{2}}+\frac{3}{8}\frac{l_{0}^{4}}{R^{4}}+...\right]+C=\frac{8\pi G}{3}\rho,$ (14)
@@
$H^{2}+\frac{k}{a^{2}}-\alpha\left(H^{2}+\frac{k}{a^{2}}\right)^{2}=\frac{8\pi G}{3}(\rho+\rho_{\Lambda}),$ (16)
where $\alpha=3l_{0}^{2}/4$, $\rho_{\Lambda}=\Lambda/(8\pi G)$. In the absence of the cosmological constant ($\Lambda=0$) and for a flat universe ($k=0$), the modified Friedmann equation is simplified as
$H^{2}-\alpha H^{4}=\frac{8\pi G}{3}\rho.$ (17)
@@
$\alpha X^{2}+X-\frac{8\pi G}{3}\rho=0,$ (18)
where we have again set $\rho_{\Lambda}=0$, for simplicity. Eq. (18) admits the following solution
$X=\frac{2}{3l_{0}^{2}}\left[1\mp\sqrt{1-8\pi Gl_{0}^{2}\rho}\right].$ (19)
@@
$H^{2}+\frac{k}{a^{2}}=\frac{8\pi G}{3}\rho\,(1+\Gamma\rho),$ (21)
where we have defined $\Gamma=2\pi Gl_{0}^{2}$.
@@
$H^{2}\approx\frac{8\pi G}{3}\Gamma\rho^{2}.$ (22)
@@ III Baryogenesis Constraints on Zero-Point Length Cosmology
Beginning with the modified first Friedmann equation
@@ IV The relation between cosmic time and temperature in zero-point length cosmology
which is the time-temperature relation in zero-point length cosmology. In the limiting case where $l_{0}\to 0$ ($\alpha=0$), the time-temperature relation simplifies to $t={1}/{2\beta T^{2}}$, which corresponds exactly to the standard general relativity result Weinberg . We note that while our modified Friedmann equations were originally derived in natural units, dimensional consistency necessitates that the right-hand side of Eq. (56) must possess dimensions of time $[t]$, and accordingly all fundamental constants have been properly restored.
@@
Figure 2: The behavior of the temperature vs time in the radiation dominated era of zero-point length cosmology for different values of the zero-point length parameter $\tilde{l}_{0}=[5\times 10^{-10},7\times 10^{-10},10^{-9}]$.
@@ V Conclusions
It is important to note that, in addition to the zero-point length correction to the area law of entropy discussed in this work, various generalization of the Bekenstein-Hawking entropy has been explored in the literatures Barrow ; Tsallis ; Kan1 ; Kan2 . While the cosmological consequences of these modified entropy (Barrow ; Tsallis ; Kan1 ; Kan2 ) on the BBN at the early stages of the Universe have been explored in Barrow2 ; Anish ; Ava1 ; Ava2 , the influences of the Kaniadakis, Tsallis and Barrow entropies on the early baryogenesis have been studied in Luciv ; Luciv2 ; Luciv3 . Therefore, it is worth constraining other modified entropies through gravitational baryogenesis. These issues are now under investigations and the results will be appeared elsewhere.
Acknowledgements.
We are grateful to Shiraz university Research Council. The work of A. Sheykhi is based upon research funded by Iran National Science Foundation (INSF) under project No. 4022705.